When 79.3 g of a particular compound is dissolved in 0.878 kg of water at 1atm pressure, the solution freezes at -1.34˚C. If the compound does not undergo ionization in solution and is nonvolatile, determine the molecular mass of the compound and freezing point depression.
[Kf = 1.858˚C/m and Tf˚ = 0˚C]
1
Expert's answer
2012-06-15T08:03:18-0400
dT=K*m*i
dT=0-(1.34)=1.34
1.34=1.858*m*1
m=1.34/1.858=0.721
m=n/m(water)
n=m*m(water)=0.721*0.878=0.633
mol
n=m/M
M=79.3 g / 0.633 mol = 125.3 g/mol